OpenStudy (anonymous):

1.Find the solution for the set of equations: a) y = x + 3 b) y = 3x + 7

OpenStudy (nirmalnema):

x+3=3x+7 2x=-4 x=-2 y=1

OpenStudy (anonymous):

@Mathbreaker

OpenStudy (anonymous):

10x i think

OpenStudy (anonymous):

as y is same for both equations, you can simply make it x+3 = 3x + 7 thus its 2x= -4 so x = -2... :)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

u r welcome